OpenStudy (anonymous):

Subtract. –52 – (–27) A. 79 B. 25 C. –25 D. –79

OpenStudy (anonymous):

Two negatives make a positive! In math, anyway.

OpenStudy (whpalmer4):

subtracting a negative number is the same as adding the opposite. 5 - (-3) = 5 + 3 = 8 NOT necessarily the case that two negatives make a positive in math, however. -3 - (-1) = -3 + 1 = -2